# Forecasting causal dynamics with universal reservoirs

### v1

@@ Abstract
An iterated multistep forecasting scheme based on recurrent neural networks (RNN) is proposed for the time series generated by causal chains with infinite memory. This forecasting strategy contains, as a particular case, the iterative prediction strategies for dynamical systems that are customary in reservoir computing. Readily computable error bounds are obtained as a function of the forecasting horizon, functional and dynamical features of the specific RNN used, and the approximation error committed by it. The framework in the paper circumvents difficult-to-verify embedding hypotheses that appear in previous references in the literature and applies to new situations like the finite-dimensional observations of functional differential equations or the deterministic parts of stochastic processes to which standard embedding techniques do not necessarily apply.
@@
22footnotetext: Faculty of Mathematics and Statistics. University of St. Gallen. Bodanstrasse 6, St. Gallen CH-9000, Switzerland. Lyudmila.Grigoryeva@unisg.ch 33footnotetext: Honorary Associate Professor at the Department of Statistics, University of Warwick, Coventry CV4 7AL, UK. Lyudmila.Grigoryeva@warwick.ac.uk 44footnotetext: Division of Mathematical Sciences, School of Physical and Mathematical Sciences, Nanyang Technological University, 21 Nanyang Link, Singapore 637371. Louw0002@e.ntu.edu.sg, Juan-Pablo.Ortega@ntu.edu.sg
@@ 1 Introduction
Recurrent neural networks (RNN) and in particular reservoir computing (RC) [Luko 09, Tana 19] have gained much visibility in recent years due to their remarkable success in the learning of the chaotic attractors of complex nonlinear infinite dimensional dynamical systems [Jaeg 04, Path 17, Path 18, Lu 18, Peco 20, Hart 20, Grig 21a, Arco 22] and in a great variety of empirical classification and forecasting applications [Wyff 10, Bute 13, Grig 14]). Even though much research has been conducted in recent years that theoretically explains this good performance, most of the work has been concentrated on the estimation [Gono 20a] and approximation [Grig 18b, Grig 18a, Gono 20b, Gono 21, Cuch 22, Gono 23a] properties of input/output systems but not in the way in which reservoir computing has been traditionally used as an RNN in order to forecast the observations of a dynamical system.
This paper aims to quantify how the universal approximation properties of a family of recurrent networks in the sense of input/output systems impact their approximation properties in the realm of the modeling and forecasting of the observations of autonomous dynamical systems and, more generally, of causal chains with infinite memory (CCIM) [Dede 07, Douk 08, Alqu 12]. More specifically, we shall quantify the forecasting performance of an RNN that has been trained to mimic the observations of a dynamical system or a causal chain to a certain accuracy as an input/output system (where the output is a time-shifted version of the system observations in the input). This result will be obtained with minimal requirements on the state-space map. The upshot of our main result is that the short-term forecasting (or path-continuation) error committed by an RNN that has been trained to uniformly approximate the dynamics of the observations up to a certain accuracy $\varepsilon>0$ is bounded above by the product of $\varepsilon$ with an exponential factor controlled by the top Lyapunov exponent of the autonomous dynamical system associated with the RNN that has been trained using the observations.
@@
  • The short-term forecasting bounds, or at least the rate at which the error grows with the forecasting horizon, are readily computable in terms of the approximation error and analytic and dynamical features of the RNN used for forecasting. The dependence on the approximation error makes it possible to use bounds developed in previous works (see [Gono 23a, Gono 23b]) that estimate it as a function of relevant architecture parameters in many popular reservoir families.
@@ 2 Preliminaries
${\bf y}_{t}=H\left(\ldots,{\bf y}_{t-2},{\bf y}_{t-1}\right)=H\left(\underline{{\bf y}_{t-1}}\right),\quad\mbox{for all $t\in\mathbb{Z}$,}$
and where the symbol $\underline{{\bf y}_{t-1}}\in({\mathbb{R}}^{d})^{\mathbb{Z}_{-}}$ denotes the semi-infinite sequence $\underline{{\bf y}_{t-1}}=\left(\ldots,{\bf y}_{t-2},{\bf y}_{t-1}\right)$. These dynamics are the deterministic counterparts of various structures with a widespread presence in time series analysis. See for instance [Dede 07, Douk 08, Alqu 12] and references therein.
@@
$U_{H}({\bf y})_{t}:=H\left(\underline{{\bf y}_{t-1}}\right),\quad\mbox{for all${\bf y}\in({\mathbb{R}}^{d})^{\mathbb{Z}}$ and $t\in\mathbb{Z}$.}$ (2.1)
@@
where $F:\mathbb{R}^{L}\times\mathbb{R}^{d}\longrightarrow\mathbb{R}^{L}$ and $h:\mathbb{R}^{L}\longrightarrow\mathbb{R}^{d}$ are a state and a readout map, respectively, $\mathbb{R}^{L}$ is, in this case, the state space and, typically, $N$ is much bigger than $d$. Some classical references in connection to this problem are [Kalm 59b, Kalm 59a, Kalm 60, Kalm 62, Baum 66, Kalm 10] or, more recently, [Sand 78, Sont 79, Dang 84, Nare 90, Matt 94, Lind 15, Grig 21b, Orte 24b]. Even though the realization problem does not always have a solution, it has been shown that if the domain of the functional $H$ is restricted to a space of uniformly bounded sequences of the type $K_{M}:=\left\{{\bf y}\in(\mathbb{R}^{d})^{\mathbb{Z}_{-}}\mid\left\|{\bf y}_{t}\right\|\leq M,\,t\in\mathbb{Z}_{-}\right\}$ for some $M>0$ and, additionally, it has a continuity condition called the fading memory property [Boyd 85, Grig 18b], then various families of state-space transformations of the type (2.2)-(2.3) can be used to approximate it uniformly. This is the case for linear systems with polynomial or neural network readouts [Boyd 85, Grig 18b, Gono 20b], state-affine systems with linear readouts [Grig 18b], or echo state networks [Grig 18a, Gono 20b, Gono 23a, Gono 21]. In other words, any of these families exhibit universality properties in the fading memory category. This allows us to conclude that given a fading memory functional $H:K_{M}\longrightarrow\mathbb{R}^{d}$ and $\varepsilon>0$, there exists a state-space system $(F,h)$ in that family to which a functional $H^{F}_{h}:K_{M}\longrightarrow\mathbb{R}^{d}$ can be associated and such that
@@
For future reference, a condition that ensures that a state-space system $(F,h)$ has a functional $H^{F}_{h}:K_{M}\longrightarrow\mathbb{R}^{d}$ associated is the so-called echo state property (ESP) [Jaeg 10]: the state equation $F$ has the ESP with respect to inputs in $K_{M}$ when for any ${\bf y}\in K_{M}$ there exists a unique $\mathbf{x}\in(\mathbb{R}^{L})^{\mathbb{Z}_{-}}$ such that (2.2) holds for any $t\in\mathbb{Z}_{-}$. In such a situation, $F$ has a filter $U^{F}:K_{M}\longrightarrow(\mathbb{R}^{L})^{\mathbb{Z}_{-}}$ associated that is uniquely determined by the relations $U^{F}({\bf y})_{t}=F\left(U^{F}({\bf y})_{t-1},{\bf y}_{t-1}\right)$. The filter $U^{F}$ induces the functional $H^{F}\colon K_{M}\longrightarrow{\mathbb{R}}^{N}$ given by $H^{F}({\bf y})=U^{F}({\bf y})_{0}.$ The functional $H^{F}_{h}:K_{M}\longrightarrow\mathbb{R}^{d}$ corresponding to the state-space system is given by $H^{F}_{h}({\bf y})=h\circ H^{F}({\bf y})$. The ESP has important dynamical implications [Grig 19, Manj 20, Manj 22]. In practice, the state space is often taken to be compact. In this case, if $F$ is a continuous map and a contraction on its first entry, the ESP is ensured. Indeed, for such a state map, it can be shown that with inputs from $\overline{B({\bf 0},M)}\subset\mathbb{R}^{d}$ there is a compact subset $X\subset\mathbb{R}^{L}$ such that $F(X\times\overline{B({\bf 0},M)})\subset X$. Stochastic generalizations of some of these statements can be found in [Manj 23, Orte 24a].
@@
$f_{(\phi,\omega,F)}\circ\phi=\Phi\circ f_{(\phi,\omega,F)}.$ (2.9)
@@
An important tool in the study of dynamical systems is the concept of ergodicity, which allows us to translate statements about space averages to averages along trajectories of the system, thus relating global analytic properties of a map to the dynamics it induces. Of particular relevance to our work is the study of nonuniform hyperbolicity and the multiplicative ergodic theorem of Oseledets [Osel 68, Ruel 79, Barr 07]. We briefly introduce the theory as applied to our context, keeping to the details and results necessary for our work. Our overview is based largely on [Barr 07, Chapter 1 and Chapter 3].
@@
$V_{i}^{-}(m):=\{v\in T_{m}M\colon\lambda^{-}(m,v)\leq\lambda_{i}^{-}(m)\},$ (2.13)
@@ 3.1 The iterated multistep forecasting scheme
for $t\in\mathbb{N},$ where we initialize $\widetilde{\bf x}_{0}=H^{F}(\underline{{\bf y}_{-1}})$ and $\widetilde{\bf y}_{0}={\bf y}_{0}.$ If we denote by ${\bf y}\mathbf{v}=\left(\ldots,{\bf y}_{-1},{\bf y}_{0},\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\right)\in({\mathbb{R}}^{d})^{\mathbb{Z}_{-}}$ the concatenation of ${\bf y}\in({\mathbb{R}}^{d})^{\mathbb{Z}_{-}}$ and $\mathbf{v}\in({\mathbb{R}}^{d})^{n}$, we can write our two sequences as
@@
We also let ${\bf x}_{t}=H^{F}(\underline{{\bf y}_{t-1}})$. While our aim is to estimate the forecasting error $\Delta{\bf y}_{t}:={\bf y}_{t}-\widetilde{\bf y}_{t},$ we do this by mapping the sequences $\underline{{\bf y}_{t}}$ and $\underline{{\bf y}_{0}}\,\underline{\widetilde{\bf y}_{t}}$ into the state-space variables ${\bf x}_{t},\widetilde{\bf x}_{t}\in\mathbb{R}^{L}$ via $H^{F}$ and studying how the error $\Delta{\bf x}_{t}:={\bf x}_{t}-\widetilde{\bf x}_{t}$ accumulates there.
@@
The reason why we called this setup simplified at the beginning of this paragraph is that in practice, one does not have at one’s disposal a full semi-infinite history of the chain that one wants to forecast but only a finite sample, say $\left\{{\bf y}_{-T},{\bf y}_{-T+1},\ldots,{\bf y}_{0}\right\}$. This fact carries in its wake two complementary sources of error that we briefly describe, together with references where they are specifically addressed:
@@
The estimation error: the universality results that we cited in the introduction guarantee the existence of uniform approximants $H_{h}^{F}$ that satisfy (3.1) in a variety of situations and for various families of state-space systems. Given a finite sample $\left\{{\bf y}_{-T},{\bf y}_{-T+1},\ldots,{\bf y}_{0}\right\}$, the approximant $H_{h}^{F}$ in the family is in general estimated using the empirical risk minimization with a prescribed loss. All estimation methods have an error associated that decreases with the sample size. Explicit bounds for this error can be found in [Gono 20a].
@@
The state starting error: as it can be seen in (3.4), the iterations that produce the forecasts are initialized using the state given by $H^{F}(\underline{{\bf y}_{-1}})\in{\mathbb{R}}^{L}$ which is uniquely determined if $F$ has the echo state property and the entire semi-infinite sequence $\underline{{\bf y}_{-1}}$ is known. When only a finite history $\left\{{\bf y}_{-T},{\bf y}_{-T+1},\ldots,{\bf y}_{0}\right\}$ is available, the initial state $H^{F}(\underline{{\bf y}_{-1}})$ can be determined only approximately, which leads to what we denominate the state starting error. If the state map $F$ is contracting in the state variable, then the corresponding filter has what is called the state forgetting property [Grig 19, Theorem 27] elsewhere known as the uniform attraction property (see [Manj 22]). This property implies that if we initialize the state equation at an arbitrary state sufficiently far in the past, then the resulting state at $t=0$ converges to $H^{F}(\underline{{\bf y}_{-1}})$ as $T\rightarrow\infty$ regardless of the initialization chosen. This means that in the presence of this property, the state $H^{F}(\underline{{\bf y}_{-1}})$ can be approximated by using a sufficiently long washout. In the dynamical systems context for which an embedding generalized synchronization exists, one can learn a map that allows what is called the cold starting of the RNN without necessarily using a washout (see [Grig 24] for a presentation of this methodology).
@@
for constants $R^{\prime}=R^{\prime}(\delta,{\bf x})$ and $R^{\prime\prime}=R^{\prime\prime}(\delta,{\bf x})$ Choosing the constant $R$ appropriately then gives us the bound in the theorem. In particular, for the first and third cases we may take $R(\delta,{\bf x}):=R^{\prime}(\delta,{\bf x})R^{\prime\prime}(\delta,{\bf x})$ and $R(\delta,{\bf x}):=R^{\prime}(\delta/4,{\bf x})R^{\prime\prime}(\delta/4,{\bf x}),$ respectively. In the second case, we note that $te^{4t\delta}$ is eventually overtaken by $e^{5t\delta},$ so there is a constant $R^{\prime\prime\prime}(\delta)$ such that $R^{\prime\prime\prime}(\delta)e^{5t\delta}\geq te^{4t\delta}$ for all $t\in{\mathbb{N}}.$ Taking $R(\delta,{\bf x}):=R^{\prime\prime\prime}(\delta/5)R^{\prime}(\delta/5,{\bf x})$ yields the desired result.
@@
$\frac{1}{C(\delta,{\bf x},i)}e^{t(\lambda_{i}({\bf x})-\delta)}\lVert{\bf v}\rVert<\lVert D\Phi^{t}({\bf x}){\bf v}\rVert<C(\delta,{\bf x},i)e^{t(\lambda_{i}({\bf x})+\delta)}\lVert{\bf v}\rVert.$ (3.9)
@@
$\lVert\left.D\Phi^{t}({\bf x})\right|_{E_{i}({\bf x})}\rVert$ $\leq C(\delta,{\bf x},i)e^{t(\lambda_{i}({\bf x})+\delta)},$ (3.10)
@@
$\lVert\left.\left(D\Phi^{t}({\bf x})\right)^{-1}\right|_{E_{i}(\widetilde{\bf x}_{t})}\rVert$ $\leq C(\delta,{\bf x},i)e^{-t(\lambda_{i}({\bf x})-\delta)}.$ (3.11)

### v2

@@ Abstract
An iterated multistep forecasting scheme based on recurrent neural networks (RNN) is proposed for the time series generated by causal chains with infinite memory. This forecasting strategy contains, as a particular case, the iterative prediction strategies for dynamical systems that are customary in reservoir computing. Explicit error bounds are obtained as a function of the forecasting horizon, functional and dynamical features of the specific RNN used, and the approximation error committed by it. In particular, the growth rate of the error is shown to be exponential and controlled by the top Lyapunov exponent of the proxy system. The framework in the paper circumvents difficult-to-verify embedding hypotheses that appear in previous references in the literature and applies to new situations like the finite-dimensional observations of functional differential equations or the deterministic parts of stochastic processes to which standard embedding techniques do not necessarily apply.
@@
22footnotetext: Mathematics and Statistics Division. University of St. Gallen. Rosenbergstrasse 22, St. Gallen CH-9000, Switzerland.Lyudmila.Grigoryeva@unisg.ch33footnotetext: Honorary Associate Professor at the Department of Statistics, University of Warwick, Coventry CV4 7AL, UK. Lyudmila.Grigoryeva@warwick.ac.uk 44footnotetext: Division of Mathematical Sciences, School of Physical and Mathematical Sciences, Nanyang Technological University, 21 Nanyang Link, Singapore 637371. Louw0002@e.ntu.edu.sg, Juan-Pablo.Ortega@ntu.edu.sg55footnotetext: Corresponding author.
@@ 1 Introduction
Recurrent neural networks (RNN) and in particular reservoir computing (RC) [Luko 09, Tana 19] have gained much visibility in recent years due to their remarkable success in the learning of the chaotic attractors of complex nonlinear infinite dimensional dynamical systems [Jaeg 04, Path 17, Path 18, Lu 18, Peco 20, Hart 20, Grig 21a, Arco 22] and in a great variety of empirical classification and forecasting applications [Wyff 10, Bute 13, Grig 14]). Even though much research has been conducted in recent years that theoretically explains this good performance, most of the work has been concentrated on the estimation [Gono 20a] and approximation [Grig 18b, Grig 18a, Gono 20b, Gono 21, Cuch 22, Gono 23] properties of input/output systems but not in the way in which reservoir computing has been traditionally used as an RNN in order to forecast the observations of a dynamical system.
This paper aims to quantify how the universal approximation properties of a family of recurrent networks in the sense of input/output systems impact their approximation properties in the realm of the modeling and forecasting of the observations of autonomous dynamical systems and, more generally, of causal chains with infinite memory (CCIM) [Dede 07, Douk 08, Alqu 12]. More specifically, we shall quantify the forecasting performance of an RNN that has been trained to mimic the observations of a dynamical system or a causal chain to a certain accuracy as an input/output system (where the output is a time-shifted version of the system observations in the input). This result will be obtained with minimal requirements on the state-space map. The upshot of our main result (Theorem 3.11) is that the short-term forecasting (or path-continuation) error committed by an RNN that has been trained to uniformly approximate the dynamics of the observations up to a certain accuracy $\varepsilon>0$ is bounded above by the product of $\varepsilon$ with an exponential factor controlled by the top Lyapunov exponent of the autonomous dynamical system associated with the RNN that has been trained using the observations.
@@
  • The rate at which the error grows with the forecasting horizon in the bounds is provided, and the dependence of the error bounds on the approximation error and analytic and dynamical features of the RNN used for forecasting is explicit. The dependence on the approximation error makes it possible to use bounds developed in previous works (see [Gono 23, Gono 24]) that estimate it as a function of relevant architecture parameters in many popular reservoir families.
Outline of the paper.
In Section 2 we give an introduction to state-space systems and causal chains with infinite memory, as well as a brief overview of the ergodic theorems for dynamical systems needed in the proofs which follow later in the paper. Section 3 develops an analysis of the error in the iterated multistep forecasting scheme, from which we derive a bound for a simplified problem in Theorem 3.1, before arriving at our main result, Theorem 3.11. Thereafter, we demonstrate how our results may be applied to the causal embedding scheme introduced in [Manj 21]. We conclude with Section 4, in which we illustrate the bound (3.12) for three different systems.
Codes and data.
All the codes and data necessary to reproduce the result in the paper are available at this GitHub repository.
@@ 2 Preliminaries
${\bf y}_{t}=H\left(\ldots,{\bf y}_{t-2},{\bf y}_{t-1}\right)=H(\underline{{\bf y}_{t-1}}),\quad\mbox{for all $t\in\mathbb{Z}$,}$
and where the symbol $\underline{{\bf y}_{t-1}}\in({\mathbb{R}}^{d})^{\mathbb{Z}_{-}}$ denotes the semi-infinite sequence $\underline{{\bf y}_{t-1}}=\left(\ldots,{\bf y}_{t-2},{\bf y}_{t-1}\right)$. These dynamics are the deterministic counterparts of various structures with a widespread presence in time series analysis. See for instance [Dede 07, Douk 08, Alqu 12] and references therein. Causal chains also have strong connections to symbolic dynamics [Das 24]. The (left) shift map $\sigma\colon\mathcal{Y}^{\mathbb{Z}}\longrightarrow\mathcal{Y}^{\mathbb{Z}},\,\sigma({\bf y})_{t}={\bf y}_{t+1}$ is related to the functional $H$ associated to the CCIM by
$H\circ\pi_{\mathbb{Z}_{-}}\circ\sigma^{t}({\bf y})=\pi_{0}\circ\sigma^{t}({\bf y})=\pi_{t}({\bf y}),\text{ for all }t\in\mathbb{Z},$
where $\pi_{\mathbb{Z}_{-}}\colon\mathcal{Y}^{\mathbb{Z}}\longrightarrow\mathcal{Y}^{\mathbb{Z}_{-}}$ is the projection onto the left infinite history, and $\pi_{t}\colon\mathcal{Y}^{\mathbb{Z}}\longrightarrow\mathcal{Y}$ the projection onto the $t$-th coordinate. Alternatively, we may write
$(\dots,{\bf y}_{t-2},{\bf y}_{t-1},H(\underline{{\bf y}_{t-1}}))=\pi_{\mathbb{Z}_{-}}\circ\sigma^{t+1}({\bf y})\text{ for all }t\in\mathbb{Z}.$
Thus the family of sequences that form the causal chain are a shift invariant subspace of $\mathcal{Y}^{\mathbb{Z}}$. While in symbolic dynamics one typically considers shift spaces over a finite alphabet, we allow the space $\mathcal{Y}$ to be more general, possibly uncountable. Typically we take $\mathcal{Y}=\mathbb{R}$. In the field of dynamical systems, symbolic dynamics supplies a very tractable means of studying integral concepts such as topological entropy, and more generally, ergodic theory [Kato 95]. A number of key examples lending insight into the properties of such quantities come from symbolic dynamics. Certain classes of dynamical systems may be approximated by shift spaces through the use of Markov partitions. The application of these ideas to general causal chains would be an interesting direction of study.
An important aspect of causal chains, which poses problems in carrying out numerical realizations of such chains, is that they depend on an infinite history of values, whereas a computer algorithm can only take as input a finite length sequence. This raises the question of the computability of causal chains. While we make some comments below on the approximation of causal chains by functions with finite length inputs, readers may look at [Lind 95, Kitc 12] for an introduction to computability theory. A further interesting reference discussing the interaction between symbolic dynamics and computability theory may be found in [Burr 22].
@@
$U_{H}({\bf y})_{t}:=H(\underline{{\bf y}_{t-1}}),\quad\mbox{for all ${\bf y}\in({\mathbb{R}}^{d})^{\mathbb{Z}}$ and $t\in\mathbb{Z}$.}$ (2.1)
@@
where $F:\mathbb{R}^{L}\times\mathbb{R}^{d}\longrightarrow\mathbb{R}^{L}$ and $h:\mathbb{R}^{L}\longrightarrow\mathbb{R}^{d}$ are a state and a readout map, respectively, $\mathbb{R}^{L}$ is, in this case, the state space and, typically, $N$ is much bigger than $d$. Some classical references in connection to this problem are [Kalm 59b, Kalm 59a, Kalm 60, Kalm 62, Baum 66, Kalm 10] or, more recently, [Sand 78, Sont 79, Dang 84, Nare 90, Matt 94, Lind 15, Grig 21b, Orte 24b]. Even though the realization problem does not always have a solution, it has been shown that if the domain of the functional $H$ is restricted to a space of uniformly bounded sequences of the type $K_{M}:=\left\{{\bf y}\in(\mathbb{R}^{d})^{\mathbb{Z}_{-}}\mid\left\|{\bf y}_{t}\right\|\leq M,\,t\in\mathbb{Z}_{-}\right\}$ for some $M>0$ and, additionally, it has a continuity condition called the fading memory property [Boyd 85, Grig 18b], then various families of state-space transformations of the type (2.2)-(2.3) can be used to approximate it uniformly. This is the case for linear systems with polynomial or neural network readouts [Boyd 85, Grig 18b, Gono 20b], state-affine systems with linear readouts [Grig 18b], or echo state networks [Grig 18a, Gono 20b, Gono 23, Gono 21]. In other words, any of these families exhibit universality properties in the fading memory category. This allows us to conclude that given a fading memory functional $H:K_{M}\longrightarrow\mathbb{R}^{d}$ and $\varepsilon>0$, there exists a state-space system $(F,h)$ in that family to which a functional $H^{F}_{h}:K_{M}\longrightarrow\mathbb{R}^{d}$ can be associated and such that
@@
For future reference, a condition that ensures that a state-space system $(F,h)$ has a functional $H^{F}_{h}:K_{M}\longrightarrow\mathbb{R}^{d}$ associated is the so-called echo state property (ESP) [Jaeg 10]: the state maps $F$ has the ESP with respect to inputs in $K_{M}$ when for any ${\bf y}\in K_{M}$ there exists a unique $\mathbf{x}\in(\mathbb{R}^{L})^{\mathbb{Z}_{-}}$ such that (2.2) holds for any $t\in\mathbb{Z}_{-}$. In such a situation, $F$ has a filter $U^{F}:K_{M}\longrightarrow(\mathbb{R}^{L})^{\mathbb{Z}_{-}}$ associated that is uniquely determined by the relations $U^{F}({\bf y})_{t}=F\left(U^{F}({\bf y})_{t-1},{\bf y}_{t-1}\right)$. The filter $U^{F}$ induces the functional $H^{F}\colon K_{M}\longrightarrow{\mathbb{R}}^{N}$ given by $H^{F}({\bf y})=U^{F}({\bf y})_{0}$. The functional $H^{F}_{h}:K_{M}\longrightarrow\mathbb{R}^{d}$ corresponding to the state-space system is given by $H^{F}_{h}({\bf y})=h\circ H^{F}({\bf y})$. The ESP has important dynamical implications [Grig 19, Manj 20, Manj 22]. In practice, the state space is often taken to be compact. In this case, if $F$ is a continuous map and a contraction on its first entry, the ESP is ensured. Indeed, for such a state map, it can be shown that with inputs from $\overline{B({\bf 0},M)}\subset\mathbb{R}^{d}$ there is a compact subset $X\subset\mathbb{R}^{L}$ such that $F(X\times\overline{B({\bf 0},M)})\subset X$. Stochastic generalizations of some of these statements can be found in [Manj 23, Orte 24a].
@@
$f_{(\phi,\omega,F)}\circ\phi=\Phi\circ f_{(\phi,\omega,F)}.$
@@
An important tool in the study of dynamical systems is the concept of ergodicity, which allows us to translate statements about space averages to averages along trajectories of the system, thus relating global analytic properties of a map to the dynamics it induces. Of particular relevance to our work is the study of nonuniform hyperbolicity and the multiplicative ergodic theorem of Oseledets [Osel 68, Ruel 79, Barr 07]. It is this theorem which forms the cornerstone of the proof, yielding the exponential factor in the final bound (3.12). We briefly introduce the theory as applied to our context, keeping to the details and results necessary for our work. Our overview is based largely on [Barr 07, Chapter 1 and Chapter 3].
@@
$V_{i}^{-}(m):=\{v\in T_{m}M\colon\lambda^{-}(m,v)\leq\lambda_{i}^{-}(m)\},$
@@ 3.1 The iterated multistep forecasting scheme
for $t\in\mathbb{N}$, and initialized with $\widetilde{\bf x}_{0}=H^{F}(\underline{{\bf y}_{-1}})$ and $\widetilde{\bf y}_{0}={\bf y}_{0}$. If we denote the concatenation of ${\bf y}\in({\mathbb{R}}^{d})^{\mathbb{Z}_{-}}$ and $\mathbf{v}\in({\mathbb{R}}^{d})^{n}$ by ${\bf y}\mathbf{v}=\left(\ldots,{\bf y}_{-1},{\bf y}_{0},\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\right)\in({\mathbb{R}}^{d})^{\mathbb{Z}_{-}}$, we can write our two sequences as
@@
We also let ${\bf x}_{t}=H^{F}(\underline{{\bf y}_{t-1}})$. While our aim is to estimate the forecasting error $\Delta{\bf y}_{t}:={\bf y}_{t}-\widetilde{\bf y}_{t}$, we do this by mapping the sequences $\underline{{\bf y}_{t}}$ and $\underline{{\bf y}_{0}}\,\underline{\widetilde{\bf y}_{t}}$ into the state-space variables ${\bf x}_{t},\widetilde{\bf x}_{t}\in\mathbb{R}^{L}$ via $H^{F}$ and studying how the error $\Delta{\bf x}_{t}:={\bf x}_{t}-\widetilde{\bf x}_{t}$ accumulates there. In Theorem 3.1, we see that the sequences separate at an exponential rate controlled by the top Lyapunov exponent of the associated dynamical system, and in Theorem 3.11, we see how this exponential separation in the state space carries over to the predictions of the time series being learnt.
@@
The reason why we called this setup simplified at the beginning of this paragraph is that, as mentioned also in the discussion of the computability of causal chains, in practice, one does not have at one’s disposal a full semi-infinite history of the chain that one wants to forecast but only a finite sample, say $\left\{{\bf y}_{-T_{0}},{\bf y}_{-T_{0}+1},\ldots,{\bf y}_{0}\right\}$. This fact carries in its wake two complementary sources of error that we briefly describe, together with references where they are specifically addressed:
@@
The estimation error: the universality results that we cited in the introduction guarantee the existence of uniform approximants $H_{h}^{F}$ that satisfy (3.1) in a variety of situations and for various families of state-space systems. Given a finite sample $\left\{{\bf y}_{-T_{0}},{\bf y}_{-T_{0}+1},\ldots,{\bf y}_{0}\right\}$, the approximant $H_{h}^{F}$ in the family is in general estimated using the empirical risk minimization with a prescribed loss. All estimation methods have an error associated that decreases with the sample size. Explicit bounds for this error can be found in [Gono 20a].
@@
The state starting error: as it can be seen in (3.4), the iterations that produce the forecasts are initialized using the state given by $H^{F}(\underline{{\bf y}_{-1}})\in{\mathbb{R}}^{L}$ which is uniquely determined if $F$ has the echo state property and the entire semi-infinite sequence $\underline{{\bf y}_{-1}}$ is known. When only a finite history $\left\{{\bf y}_{-T_{0}},{\bf y}_{-T_{0}+1},\ldots,{\bf y}_{0}\right\}$ is available, the initial state $H^{F}(\underline{{\bf y}_{-1}})$ can be determined only approximately, which leads to what we denominate the state starting error. If the state map $F$ is contracting in the state variable, then the corresponding filter has what is called the state forgetting property [Grig 19, Theorem 27] elsewhere known as the uniform attraction property (see [Manj 22]). This property implies that if we initialize the state equation at an arbitrary state sufficiently far in the past, then the resulting state at $t=0$ converges to $H^{F}(\underline{{\bf y}_{-1}})$ as $T_{0}\rightarrow\infty$ regardless of the initialization chosen. This means that in the presence of this property, the state $H^{F}(\underline{{\bf y}_{-1}})$ can be approximated by using a sufficiently long washout. In the dynamical systems context for which an embedding generalized synchronization exists, one can learn a map that allows what is called the cold starting of the RNN without necessarily using a washout (see [Grig 24] for a presentation of this methodology).
@@
for constants $R^{\prime}=R^{\prime}(\delta,{\bf x})$ and $R^{\prime\prime}=R^{\prime\prime}(\delta,{\bf x})$. Choosing the constant $R$ appropriately then gives us the bound in the theorem. In particular, for the first and third cases we may take $R(\delta,{\bf x}):=R^{\prime}(\delta,{\bf x})R^{\prime\prime}(\delta,{\bf x})$ and $R(\delta,{\bf x}):=R^{\prime}(\delta/4,{\bf x})R^{\prime\prime}(\delta/4,{\bf x})$, respectively. We note that the second case occurs only when $\lambda_{1}>0$; in that case, we can ensure by shrinking $\delta$ that we are always in the first case and not in the second one.
@@
$\frac{1}{C(\delta,{\bf x},i)}e^{t(\lambda_{i}({\bf x})-\delta)}\lVert{\bf v}\rVert<\lVert D\Phi^{t}({\bf x}){\bf v}\rVert<C(\delta,{\bf x},i)e^{t(\lambda_{i}({\bf x})+\delta)}\lVert{\bf v}\rVert.$
@@
$\lVert\left.D\Phi^{t}({\bf x})\right|_{E_{i}({\bf x})}\rVert$ $\leq C(\delta,{\bf x},i)e^{t(\lambda_{i}({\bf x})+\delta)},$
@@
$\lVert\left.\left(D\Phi^{t}({\bf x})\right)^{-1}\right|_{E_{i}(\widetilde{\bf x}_{t})}\rVert$ $\leq C(\delta,{\bf x},i)e^{-t(\lambda_{i}({\bf x})-\delta)}.$